OpenStudy (anonymous):

The town of Goodland, Kansas, claims that it has one of the world's largest easels. It holds an enlargement of a van Gogh painting that is 24 feet wide. The original painting is 58 cm wide and 73 cm tall. If the reproduction is simlar to the original, what is the height of the reproduction to the nearest foot? A.28 feet B.23 feet C.27 feet D.30 feet

OpenStudy (anonymous):

ratio of original is 58/73 new one is 24/x 58/73 = 24/x, solve for x

OpenStudy (anonymous):

x = 24/(58/73) = 30.2 D. 30

OpenStudy (anonymous):

What is the similarity ratio of these two parallelograms?